OpenStudy (anonymous):

if B is the midpoint of the line segment AC, AB=5x, and BC=3x+4 what is BC?

OpenStudy (anonymous):

by definition of mid point, Length AB is equal to Length BC.

OpenStudy (anonymous):

We know that AB = 5x, and BC = 3x+4, and we know AB==BC, so... 5x=3x+4.

OpenStudy (anonymous):

you said : if B is the midpoint of the line segment AC, AB=5x, and BC=3x+4 what is BC?

OpenStudy (anonymous):

You dont have AC. ..

OpenStudy (anonymous):

I want you to solve 5x=3x+4. get x, then we know that BC (what you are looking for) is 3x+4.

OpenStudy (anonymous):

You were NOT given the length of AC, correct?

OpenStudy (anonymous):

AB and BC meet at point B right? So B is the halfway of your line segment. First: set AB and BC = to each other: 5x=3x+4 Second: solve for x. Third: go back and plug in the number you got for x

OpenStudy (anonymous):

5x=3x+4 move over your 3x 5x-3x=4 so: 2x=4 x=2

OpenStudy (anonymous):

AB=5x so AB=5 * 2= 10 BC=3x+4 So BC= (3*2) +4 = 6+4= 10

OpenStudy (anonymous):

AB and BC are = because B is the mid point of AC So AC= 20

OpenStudy (anonymous):

You're very welcome. My Math teacher the past 2 years was kinda difficult to learn from so I helped my friends a lot